OpenStudy (anonymous):

can someone explain to me how to solve the following inequality for x please i will medal and fan

OpenStudy (anonymous):

\[-6 + 8\ge (-136)-14x\]

OpenStudy (anonymous):

i mean -6x

OpenStudy (anonymous):

\[-6x+8\ge-136-14x\]

OpenStudy (anonymous):

Well first you have to get x one one side

OpenStudy (anonymous):

The final answer is \[x \ge -18\]